OpenStudy (anonymous):

How do I find the zeros of 2x^2+2x-4

OpenStudy (anonymous):

look at the signs in front of each term so +2x^2, +2x, and -4. + to +, there is no change, but from + to - there is a change, so 1 zero

OpenStudy (anonymous):

graph it

OpenStudy (anonymous):

factor it and find zeros.

OpenStudy (jdoe0001):

\(\large {2x^2+2x-4\implies 2x^2+2x-4=0\implies 2({\color{brown}{ x^2+x-2}})=0 \\ \quad \\ \begin{array}{llll} {\color{brown}{ x^2\quad +x\quad -2}}\\ \qquad \quad \uparrow \qquad \uparrow \\ \qquad 2-1\quad 2\cdot -1 \end{array} }\)